OpenStudy (anonymous):

solve each system by substitution. 4x-2y=-14 -12x+6y=42

hero (hero):

Hint: Divide the second equation by 3

OpenStudy (anonymous):

i have y = 2x - 7 and -6x+3=21

hero (hero):

You should have divided the second equation by 2. Then you would have had the following: 4x -2y = -14 -4x + 2y = 14 Then it would have been obvious that the addition of both equations lead to a 0 = 0 scenario